OpenStudy (anonymous):

Solve using any method. x² + 4x - 32 = 0

OpenStudy (anonymous):

It factors! Watch...

OpenStudy (anonymous):

what two factor of 32 will add or subtract to give you 4

OpenStudy (anonymous):

-32=8*(-4)

OpenStudy (anonymous):

so \[x ^{2}+4x-32=0\] can be changed into (x+8)*(x-4)=0

OpenStudy (anonymous):

my method is to watch what everyone gets...:)

OpenStudy (anonymous):

32 = 2 * 2 * 2 * 2 * 2 Seeing as there's a - in the last term, that means you're going to have a + and a - when you factor... 2*2*2 - 2*2 8 - 4 4 \[(x-4) (x+8) = 0\] solve each one now \[x-4=0 \rightarrow x=4\] \[x+8=0 \rightarrow x=-8\]